### v2

# Leveraging Local Patch Alignment to Seam-cutting for Large Parallax Image Stitching

Abstract
Seam cutting methods have been proven effective in the composition step of image stitching, especially for images with parallax. However, current seam cutting can be seen as the subsequent step after the image alignment is settled. Its effectiveness usually depends on the fact that images can be roughly aligned such that a local region exists where an unnoticeable seam can be found. Current alignment methods often fall short of expectations for images with large parallax, and most efforts are devoted to improving the alignment accuracy.
In this paper, we argue that by adding a simple Local Patch Alignment Module (LPAM) into the seam cutting, the final result can be efficiently improved for large parallax image stitching. Concretely, we first evaluate the quality of pixels along the estimated seam of the seam cutting method. Then, for pixels with low qualities, we separate their enclosing patches in the aligned images and locally align them by constructing modified dense correspondences via SIFT flow. Finally, we composite the aligned patches via seam cutting and merge them into the original aligned result to generate the final mosaic. Experiments show that introducing LPAM can effectively and efficiently improve the stitching results.
Index Terms:
Image stitching, large parallax, seam-cutting, local patch alignment, SIFT flow.
I Introduction
Image stitching has been well-studied for decades in computer vision [1], which mainly involves image alignment [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], composition and blending [15, 16, 17, 18, 19, 20, 21]. In consumer-level photography, parallax is the most intractable issue due to the unconstrained shooting environment. Existing alignment methods may still fail to handle it. Therefore, it is usually the image composition that determines the quality of the final stitching result.
seam cutting [18, 17, 19, 22, 20, 21] and seam-driven approaches [23, 24, 25], which aim to find an optimal seam in the overlapping region from one or finite alignment candidates, have been proven effective in handing parallax. However, as a second step after image alignment is settled in image stitching, the effectiveness highly depends on the alignment accuracy. It means that the images should at least be roughly aligned such that there exists a local region where an unnoticeable seam can be found. For images with large parallax, existing alignment methods may fail to align them locally (see Fig. 1(b)), resulting in the seam cutting failing to produce visually pleasing results (see Fig. 1(c,d)).
(a) Input images
(b) Averaged aligned images
(c) Method [19]
(d) Method [20]
(e) Our method
Figure 1: Issues of the state-of-the-art seam cutting methods for large parallax image stitching (marked by red arrows). (b) Aligned images are obtained via method [6], where the overlapping region is averaged. The curbs are totally misaligned (Best view in color and zoom in).
In this paper, we argue that by adding a simple Local Patch Alignment Module (LPAM) into the seam cutting, the final result can be efficiently improved for large parallax image stitching. Specifically, we first introduce a seam quality evaluation algorithm to evaluate the quality of pixels along the seam and identify the “problematic” pixels. Then, for these pixels, we separate their enclosing patches in the aligned images and locally align them by constructing modified dense correspondences via SIFT flow. Finally, we composite the aligned patches via seam cutting and merge them into the former aligned result to generate the final mosaic. Experiments show that adding our LPAM into the seam cutting can effectively and efficiently improve the stitching results (see Fig. 1(e)).
Our method has two major contributions.
  • We propose a local alignment and seam cutting method that can improve the performance of the large parallax image stitching methods.
  • Our method is a plug-and-play module that can be effectively and efficiently applied to any seam cutting method.
II Related Work
We briefly review the approaches most related to our work. Comprehensive reviews of image stitching algorithms are given in [1, 26, 27].
II-A Seam cutting Methods
The seam cutting approach is a powerful and widely used composition method [15, 28, 16, 17, 18, 19, 29], which is usually expressed in terms of energy minimization and minimize it via graph-cut optimization [30]. Agarwala et al. [28] applied seam cutting to generate “digital photomontage”, they defined the energy function as the sum of a data penalty selecting by the user and an interaction penalty involving the colors, gradients, and edges of an image.
Zhang et al. [16] combined alignment error and color difference in the energy function to find a better seam that avoids splitting misaligned pixels with similar colors. Li and Liao [18] considered the nonlinearity and nonuniformity of human perception into seam cutting and defined a sigmoid-metric color difference and saliency weights in the energy function. To avoid objects from being cropped, omitted, or duplicated, Herrmann et al. [17] utilized object detection to match objects between images, and introduced three new terms to the traditional energy function that address the above issues, respectively. Liao et al. [19] indicated that the seam with minimal cost is not necessarily optimal in terms of human perception. Then, they proposed an iterative seam estimation method where a quality evaluation algorithm for pixels along the seam is designed to guide the iteration procedure. Dai et al. [22] proposed the first end-to-end deep learning framework for image composition. They cast the seam cutting as an image-blending problem and designed an edge-guided network to regress the blending weights and seamlessly produce the stitched image. Li and Zhou [20] utilize the quaternion representation of the color image to define the energy function and perform seam cutting completely in the quaternion domain.
II-B Seam-driven Methods
The seam-driven strategy was first proposed by Gao et al. [23]. They indicated that the perceptually best result is not necessarily from the best global alignment. Thus, they generated finite alignment candidates from multiple RANSAC [31] procedures and applied seam cutting to estimate multiple seams from these candidates. Then, a seam quality metric is defined to evaluate these seams, and the seam with minimal cost is chosen to produce a final mosaic. Zhang and Liu [24] improved the screening process of alignment candidates by combining homography with content-preserving warps. They used the seam cost of the energy function to evaluate the seam quality. Lin et al. [25] generated the alignment candidates via a superpixel-based feature grouping and a structure-preserving warp. The warp is gradually improved via adaptive feature weighting. They also defined a seam quality metric based on the zero-mean normalized cross-correlation (ZNCC) score. Zhang et al. [21] proposed an iterative seam-based stitching method that leverages dense flow estimation to directly estimate seam in the input images and use it to solve a spatial smooth warp.
In both seam cutting and seam-driven methods, their main focus is defining various energy functions to find an optimal seam from the settled pre-aligned images. For images with large parallax, these methods may easily fail if the alignment methods fail to locally align the images. Our method, on the other hand, improves the performance by realigning the local misaligned patches where the estimated seam passes.
Figure 2: An overview of the proposed local patch alignment module.
III Proposed architecture
As depicted in Fig. 2, our architecture consists of two stages, seam quality evaluation and patch extraction, succeeded by patch alignment and seam merging. Subsequently, we first provide a preliminary of how seam cutting works, and then present a detailed description of the two stages.
III-A Preliminary
Given aligned target image $I_{0}$ and reference image $I_{1}$, for the overlapping region $\mathbf{O}$, seam cutting method estimates an initial seam $S_{0}$ by finding a labeling $l$ that minimizes the energy function
$E(l)=\sum_{p\in\mathbf{O}}D_{p}(l_{p})+\sum_{(p,q)\in N}S_{p,q}(l_{p},l_{q})$ (1)
where the data penalty $D_{p}$ measures the cost of assigning a label $l_{p}$ to pixel $p$, the smoothness penalty $S_{p,q}$ measures the cost of assigning a pair of labels $(l_{p},l_{q})$ to a pair of neighboring pixels $(p,q)\in N$. Eq. (1) can be minimized via graph-cut optimization [30].
III-B Seam quality evaluation and patch extraction
Since the seam cutting methods mainly fail due to the structure misalignment, we use the structural similarity (SSIM) index [32] to evaluate each pixel $p_{i}$ on the seam $S_{0}$. Concretely, we extract two local patches $P_{0}$ and $P_{1}$ centered at $p_{i}$, and define the evaluation error for $p_{i}$ as
$Q(p_{i})=1-\mathrm{SSIM}(P_{0}(p_{i}),P_{1}(p_{i})),$ (2)
Fig. 2 illustrates one example of the seam quality evaluation for seam $S_{0}$, pixels with high errors usually mean misaligned parts.
Due to the advanced alignment methods, even for images with large parallax, the seam only experiences partial failure. Our goal is to “repair” these failed parts. As noted in [19], a plausible seam should have a relatively small error $Q$ for every pixel. Thus, for pixels with high errors, we separate them in the seam for subsequent alignment as follows,
  • if $\max(Q)\leq k\cdot\mathrm{mean}(Q)$, we consider $S_{0}$ as a plausible seam;
  • else, we calculate a threshold $\tau$ via Ostu’s method [33], and consider $p_{i}$ with $Q(p_{i})\geq\tau$ as misaligned.
We set $k$ to 1.5 in our method. The threshold $\tau$ divides the seam into multiple components. Notably, the threshold might not be sufficiently precise to identify all misaligned pixels. However, we allow for some some false positives because the subsequent alignment process does not yield inferior results. For each misaligned component $\mathbf{C}_{i}$, we separate its enclosing rectangular patches $\mathbf{P}^{i}_{0}$ and $\mathbf{P}^{i}_{1}$ from the corresponding aligned images (see Fig. 2).
III-C Patch Alignment and seam merging
Figure 3: Illustration for patch alignment and seam merging process. Artifacts are denoted by red arrows. From left to right in (a,b,c,d): vector flow, local stitching seam, and local composited result. For clarity, only part of the vector flow is presented. From (a) to (d): results of original $V^{i}$, modified $f(V^{i})$ with $\beta=1,4,8$ in Eq. (5), respectively. The green rectangle in the middle of each subfigure indicates the position of the separated patch and the merging of the local seam. The right image in each subfigure only shows the simple composited results without gradient domain fusion (Best view in color and zoom in).
The separated patches exhibit structure misalignment, we assume that there is either no or only a small parallax between these patches. Thus, we can utilize SIFT flow [34] to construct dense correspondences among the pixels within the patches. Theoretically, for each pixel $p_{j}$ in $\mathbf{P}_{0}^{i}$, there exists a shift vector $V^{i}(p_{j})$ such that
$\|\mathbf{P}_{0}^{i}(p_{j}+V^{i}(p_{j}))-\mathbf{P}_{1}^{i}(p_{j})\|=0$ (3)
Then, the two patches can be aligned via the vector flow $V^{i}$. However, if the boundary pixels of the patch $\mathbf{P}_{0}^{i}$ have large shift vectors, merging the aligned patches into the pre-aligned results may introduce more artifacts around the patch, as shown in Fig. 3. To address this, we smoothly modify the vector flow and realign the two patches as
$\tilde{\mathbf{P}}_{0}^{i}(p_{j})=\mathbf{P}_{0}^{i}(p_{j}+f(V^{i}(p_{j}))),\quad\tilde{\mathbf{P}}_{1}^{i}(p_{j})=\mathbf{P}_{1}^{i}(p_{j})$ (4)
where $\tilde{\mathbf{P}}_{0}^{i}$ ($\tilde{\mathbf{P}}_{1}^{i}$) is the warped target (reference) patch, $f$ is the sigmoid function
$f(t)=\frac{1}{1+e^{-\beta(t-0.5)}}$ (5)
For horizontal patch alignment, $t$ represents the $x$-coordinate of a pixel within patch $\mathbf{P}_{0}^{i}$, normalized to $[0,1]$. $\beta$ controls the rate of change of the function $f$. The sigmoid function ensures that the aligned patches are smoothly consistent with $I_{0}$ when $t=0$ and with $I_{1}$ when $t=1$. As shown in Fig. 3, the modified vector flow under different settings of $\beta$ is presented. In our method, $\beta$ is set to 8.
Due to the vector flow modification, the separate patches are only partially aligned. Thus, the seam cutting is still required to composite these patches. Note that the local seam $s_{i}$ in the aligned patches must be connected with the initial seam $S_{0}$ end to end. Otherwise, artifacts may occur at the junction of the two seams. To prevent introducing new artifacts, we constrain the local seam to connect with the initial seam by finding a label $l^{i}$ in which the data penalty for pixel $p_{j}\in\mathbf{P}_{0}^{i}$ is
  • $D_{p_{j}}(0)=0,D_{p_{j}}(1)=\infty$, if $p_{j}\in\partial\mathbf{P}_{0}^{i}$ with $l_{p_{j}}=0$ in the initial seam $S_{0}$;
  • $D_{p_{j}}(0)=\infty,D_{p_{j}}(1)=0$, if $p_{j}\in\partial\mathbf{P}_{0}^{i}$ with $l_{p_{j}}=1$ in the initial seam $S_{0}$;
  • $D_{p_{j}}(0)=0,D_{p_{j}}(1)=0$, otherwise,
where $\infty$ is set to avoid mislabeling, $\partial\mathbf{P}_{0}^{i}$ represents the border of $\mathbf{P}_{0}^{i}$. Fig. 3 shows an example including the seam merging process. We can see that the local seam is accurately merged with the initial seam.
We summarize our method in Algorithm 1.
0:  two aligned images $I_{0}$, $I_{1}$ and an initial seam $S_{0}$;
0:  final stitching seam $S_{*}$;
1:  evaluate the pixels along the seam via seam quality evaluation in Eq. (2);
2:  identify whether there are misaligned pixels along the seam;
3:  if misaligned pixels exist then
4:     extract misaligned components $\{\mathbf{C}_{i}\}_{i=1}^{n}$ in seam $S_{0}$ based on threshold $\tau$;
5:     for each $\mathbf{C}_{i}$ do
6:        separate the enclosing patches $\mathbf{P}^{i}_{0}$ and $\mathbf{P}^{i}_{1}$ from $I_{0}$ and $I_{1}$, respectively;
7:        estimate the SIFT flow [34] between $\mathbf{P}^{i}_{0}$ and $\mathbf{P}^{i}_{1}$ and obtain the vector flow $V_{i}$;
8:        modify the vector flow via the sigmoid function $f$ in Eq. (5);
9:        align $\mathbf{P}^{i}_{0}$ and $\mathbf{P}^{i}_{1}$ via $f(V_{i})$;
10:        calculate a local stitching seam $s_{i}$ in the two aligned patches;
11:        obtain a new seam $S_{i}$ by merging $s_{i}$ into the last seam in the corresponding position.
12:     end for
13:  end if
14:  return  $S_{*}=S_{n}$
Algorithm 1 Local patch alignment module for large parallax image stitching.
IV Experimental evaluations
Figure 4: Our image dataset for experimental evaluation.
In our experiments, the patch size for seam quality evaluation is set to $21\times 21$. Input images are aligned using REW [6], a state-of-the-art warping method. To comprehensively assess the proposed method, LPAM is applied to several advanced seam cutting methods, including basic Euclidian-based seam cutting (Baseline), perception-based seam cutting (Perception) [18], iterative seam cutting (Iterative) [19], quaternion-based seam cutting (Qua-ternion) [20], the final mosaic is generated by applying gradient domain fusion [35] to the seam cutting result. We test our method on three large parallax image datasets: 35 image pairs from Parallax [24], 24 image pairs from SEAGULL [25], and 20 image pairs from our dataset, as shown in Fig. 4. Our dataset is available in our code [^1] https://github.com/tlliao/LPAM_seam-cutting.
IV-A Quantitative evaluation
We introduce four classical metrics, RMSE, PSNR, SSIM [32], and ZNCC [25] to evaluate the quality of the initial seam and final seam after adding the LPAM. For the stitching seam $S$, the quality errors are defined as
$\mathrm{RMSE}(S)=\frac{1}{N}\sum_{i=1}^{N}\mathrm{RMSE}(P_{0}(p_{i}),P_{1}(p_{i}))$ (6)
$\mathrm{PSNR}(S)=\frac{1}{N}\sum_{i=1}^{N}\mathrm{PSNR}(P_{0}(p_{i}),P_{1}(p_{i}))$ (7)
$\mathrm{SSIM}(S)=\frac{1}{N}\sum_{i=1}^{N}\mathrm{SSIM}(P_{0}(p_{i}),P_{1}(p_{i}))$ (8)
$\mathrm{ZNCC}(S)=\frac{1}{N}\sum_{i=1}^{N}\frac{1-\mathrm{ZNCC}(P_{0}(p_{i}),P_{1}(p_{i}))}{2}$ (9)
where $p_{i}$ is the pixel on the seam, $N$ is the number of the pixels on the seam, $P_{0}$, $P_{1}$ is the patch centered at $p_{i}$ in $I_{0}$, $I_{1}$ respectively.
Table I presents the quantitative evaluation results of different seam cutting methods before and after adding LPAM. Evidently, LPAM can substantially improve the seam quality of any seam cutting method across all large parallax image datasets. We also show some of the seam comparison results in Fig. 5, where the seams are shown using colormaps based on the quality error defined by Eq. (2). Through local alignment of misaligned patches, our method effectively improves the quality of seam cutting and the final visual stitching results.
Figure 5: Visual seam quality comparisons w/ and w/o LPAM. Top: initial seams w/o LPAM. Bottom: final seams w/ LPAM.
Our method improves the seam cutting by locally realigning the misaligned patches along the stitching seam. To verify its effectiveness, we also compare the global alignment accuracy of the overlapping region in different seam cutting methods, both before and after incorporating LPAM. The comparison results are shown in Table II. Low PSNR and SSIM values indicate that accurately aligning the input images is challenging. Our LPAM can continuously improve the alignment accuracy, regardless of the estimated stitching seam, thus generating visually superior stitching results.
TABLE I: Quantitative evaluation of seam cutting w/ and w/o LPAM on different image datasets.
Dataset Parallax [24] SEAGULL [25] Our dataset
RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ ZNCC $\downarrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ ZNCC $\downarrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ ZNCC $\downarrow$
Baseline 0.088 31.01 0.750 0.119 0.106 26.20 0.638 0.163 0.079 28.50 0.756 0.114
Baseline+LPAM 0.078 31.90 0.784 0.099 0.096 28.31 0.696 0.138 0.075 28.78 0.780 0.103
Perception [18] 0.084 30.03 0.749 0.102 0.114 24.43 0.613 0.159 0.081 28.11 0.756 0.106
Perception+LPAM 0.074 31.23 0.795 0.084 0.103 25.37 0.682 0.127 0.072 28.82 0.801 0.084
Iterative [19] 0.086 29.83 0.762 0.107 0.114 24.43 0.613 0.159 0.081 28.46 0.761 0.113
Iterative+LPAM 0.076 30.87 0.800 0.087 0.103 25.37 0.682 0.127 0.072 28.91 0.786 0.101
Quaternion [20] 0.089 28.70 0.746 0.099 0.116 25.00 0.613 0.167 0.073 28.50 0.778 0.089
Quaternion+LPAM 0.080 29.41 0.783 0.085 0.105 26.09 0.677 0.136 0.067 29.25 0.812 0.078
TABLE II: Quantitative evaluation of global alignment accuracy w/ and w/o LPAM on different image datasets.
Dataset Parallax [24] SEAGULL [25] Our dataset
PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
REW [6] 13.539 0.630 13.669 0.551 14.928 0.682
Baseline+LPAM 13.591 0.636 13.739 0.560 14.966 0.685
Perception+LPAM 13.589 0.636 13.758 0.564 14.973 0.687
Iterative+LPAM 13.593 0.636 13.759 0.561 14.974 0.685
Quaternion+LPAM 13.602 0.636 13.739 0.561 15.009 0.689
IV-B Visual Comparison
We comprehensively compare the visual stitching results of our method with those of methods integrating the seam cutting technique. These include Microsoft ICE [^2] https://www.microsoft.com/en-us/research/project/image-composite-editor, [18, 19, 20], Parallax [24], and SEAGULL [25]. Fig. 6 shows a particularly challenging example sourced from [24]. The input images and the result of Parallax are retrieved from the website [^3] http://web.cecs.pdx.edu/~fliu/project/stitch/. Notably, Microsoft ICE, the baseline method, and Iterative seam cutting [19] fail to produce satisfactory results (obvious artifacts are marked by green arrows). The simple integration of our LPAM improves the stitching quality and achieves visual results comparable to those of Parallax and SEAGULL. It should be noted that Parallax and SEAGULL need to estimate multiple seam cutting results from multiple alignment candidates and then select the optimal one. In contrast, our LPAM provides a more straightforward and effective solution for improving stitching quality.
(a) Input images
(b) Microsoft ICE
(c) Parallax [24]
(d) SEAGULL [25]
(e) Baseline
(f) Baseline+LPAM
(g) Perception [18]
(h) Perception+LPAM
(i) Iterative [19]
(j) Iterative+LPAM
(k) Quaternion [20]
(l) Quaternion+LPAM
Figure 6: Comparisons of image stitching results via seam cutting methods.
IV-C Computational Efficiency
We evaluate the computational efficiency of LPAM when integrated with other seam cutting methods. Experiments are conducted on three image datasets using a 3.6GHz Intel Core i7 CPU with 16GB memory. We record the elapsed time for each seam cutting method [^4] Time recorded is solely for seam cutting, excluding image alignment. and for LPAM individually. The results are presented in Table III. In LPAM, constructing dense correspondence dominates the elapsed time. This process is influenced by the number and resolution of extracted patches, which vary significantly across different image pairs and datasets. Notably, when applied to various seam cutting methods, LPAM’s processing time remains relatively stable, ranging from 6-10 seconds [^5] LPAM is currently implemented in MATLAB, integrating with C++ could further improve its efficiency..
TABLE III: Elapsed time of applying LPAM to different seam cutting methods on the three image datasets.
Dataset Parallax [24] SEAGULL [25] Our dataset
Baseline 7.54 4.11 5.26
+LPAM 7.77 7.34 6.89
Perception [18] 9.21 3.61 6.54
+LPAM 8.96 10.12 7.90
Iterative [19] 112.04 201.06 153.07
+LPAM 8.85 9.15 7.47
Quaternion [20] 8.95 3.51 9.59
+LPAM 9.25 10.17 9.57
(a) Input images
(b) Initial stitching seam and result
(c) Stitching seam and result w/o LPAM
(d) Stitching seam and result w/ LPAM
Figure 7: Failure examples of our method (Best view in color and zoom in).
IV-D Discussion and Limitation
Our method assumes that the stitching seam experiences partial failure such that only several patches along the seam need realignment. In our implementation, a seam is considered to be plausible if $\max(Q)\leq 1.5\cdot\mathrm{mean}(Q)$. Nevertheless, as shown in Fig. 7(a,b), this inequality can still hold even if most pixels along the seam fail. In such cases, our method fails since it regards the seam as plausible, and no patches are extracted for realignment. Designing a superior algorithm to identify the “problematic” pixels along the stitching seam could potentially resolve this issue. We leave this as future work.
Furthermore, in Eqs. (4,5), we introduce a modification to the vector flow between separated patches to smoothly merge them into pre-aligned results. However, when significant structural misalignment exists within small patches, this modification is insufficient to ensure smoothness, as shown in Fig. 7(c,d). The extracted patch (marked by the black box) is too small for effective smoothness modification. Consequently, our LPAM struggles to strike a balance between accurate alignment and smooth merging. Enlarging the patch size could yield smoother merging. We leave this as future work.
V Conclusion
In this paper, we propose a simple local patch alignment module that can effectively and efficiently improve the stitching performance of the seam cutting methods for large parallax image stitching. Firstly, we evaluate the pixels along the initial seam and extract misaligned components of the seam. For each such component, we separate the enclosing rectangular patches in the pre-aligned images and construct modified dense correspondence using SIFT flow to align these patches. Finally, we calculate a local seam and merge it into the initial seam. Experimental results on various stitching datasets demonstrate the superiority of our method.
Acknowledgment
This work is partially supported by the Natural Science Foundation of Henan Province under Grant 222300420140.